A person inhales air richer in O2 and exhales air richer in CO2 and water vapor. During each hour of sleep, a person exhales a total of about 300. L of this CO2−enriched and

H2O-enriched air.

(a) If the partial pressures of CO2 and H2O in exhaled air are each 29.3 torr at 37.0 C, calculate the masses of CO2 and of H2O exhaled in 1 h of sleep

____g CO2

____g H2O

(b) How many grams of body mass does the person lose in 8 h sleep if all the CO2 and H2O exhaled come from the metabolism of glucose? Assume that the mass lost is equal to the mass of glucose that is converted to CO2 and H2O.

C6H12O6(s) + 6 O2(g) → 6 CO2(g) + 6 H2O(g)

___g
Answer #1

a) Using ideal gas equation

PV = nRT

P = 29.3 torr = 29.3/760 atm

V = 300 L

Temperature = 273 + 37.0 = 310.0 K

29.3/760 * 300 = n * 0.0821 * 310.0

n = 0.4544 moles

Mass of CO2 = number of moles * molar mass = 0.4544 mol * 44 g/mol = 19.99 grams

Mass of H2O = number of moles * molar mass = 0.4544 mol * 18 g/mol = 8.18 grams

(b)

As per the balanced reaction, 1 mole of glucose will form 6 moles of CO2 and 6 moles of H2O

Moles of glucose lost = (0.4544)/6 = 0.075733 moles

Molar mass of glucose (C6H12O6) = 180.15 g/mol

Mass lost = number of moles * molar mass = 0.075733 mol * 180.15 g/mol = 13.64 grams

Note - Post any doubts/queries in comments section.